If a 25 ppm Nitrogen dioxide in the air contains 25 mg of the compound in every liter of solution. How many milligrams are contained in 2.5 L sample of the air?


1
Expert's answer
2022-02-23T15:39:02-0500

ppm = (milligrams of solute) / (liters of solution).


25=mg/2.5

So mg=25×2.5=62.5milligrams